A surveyor is "shooting a line" to a point on a tree 55m from his current position. After rotating his survying equipment 25* to the left he shoots another line to a point on a fence post 25m away
a) determine distance between the point on the tree to the point on the fence post. Shoe appopriate formula,substutions, and work. Give the distance to the nearest hundreth of a meter.
b) Use the laws of Sine to find thwe measure of c) Find the m
Answer by RAY100(1637)

Using Law of Cosines on initial info
.
c^2 = a^2 +b^2 -2abcosc
.
c^2 = 25^2 +55^2 -2 (25)(55) cos 25
.
c^2 = 1157.654
.
c= 34.0243=34.02 m
.
Using the law of sines
.
34.02/sin25= 80.5084 = 25/sina = 55/sinb
.
sin a = .3105,,,,a = 18.09 deg
.
sin b = .683,,,,,a= 43 or 136.965 deg,,,,,,,use 137 deg
.
Checking
.
a +b +c = 18 +137 +25 = 180,,,,,ok